# Quantifying uncertainty in climate projections with conformal ensembles

### v1

@@ 4.5 Sensitivity Analysis
Figure 4 shows the out-of-sample test SW distances for EA and GP across 10 different calibration size settings (50, 100, …, 450) and each climatic variable TAS, TMAX, and PR. Each line represents the median SW distance across all perfect model experiments, and the uncertainty bands represent 2 standard errors (2SE) from the median. As expected, UQ skill shows a “U” shaped performance curve where small and large calibration sets result in poor test UQ. For the GP method, this result is natural, considering a larger calibration set leaves less data for training. We included the EA method to show that this effect is not entirely due to insufficient training data. Because the EA method has no training step, its similarly shaped performance curve indicates that larger calibration sets might include residual functions that are irrelevant for future prediction. Again, this result is somewhat expected because the distributions of TAS, TMAX, and PR evolve over time. By including residuals too far in the past, our projection set might contain information that is irrelevant or even counter to future projections. Thus, even if sample size was not a factor, for non-stationary climatic variables, we may still need to limit the size or time period of the calibration set.
@@ 5.2 Forced response in the tails
Given the positive results in Section 5.1, confirming that CE provides reliable UQ for reanalysis data, we use our method to project TAS and PR fields over the next 30 years (May 2024 - Dec. 2054). We use the GP model as our ensemble analysis function, trained on Jan. 1940 - Nov. 2015 data for TAS and Jan 1940 - Nov. 1990 data for PR, and construct CEs on Jul. 2007 - Mar. 2024 data for TAS and Dec 1990 - Mar. 2024 data for PR. Unlike in Section 5.1, we can use the full calibration set to construct each CE. Using the GP, we make three projections over all months between 2024 and 2054 using the IMV, IMV(BC), and the CE to quantify uncertainty and denote these combinations as GP + IMV, GP + IMV(BC), and GP + CE, respectively.

### v2

@@ 4.5 Sensitivity Analysis
Figure 4 shows the out-of-sample test SW distances for EA and GP across 10 different calibration size settings ($n_{2}$ = 50, 100, …, 450) and each climatic variable TAS, TMAX, and PR. Each line represents the median SW distance across all perfect model experiments, and the uncertainty bands represent 2 standard errors (2SE) from the median. As expected, UQ skill shows a “U” shaped performance curve where small and large calibration sets result in poor test UQ. For the GP method, this result is natural, considering a larger calibration set leaves less data for training. We included the EA method to show that this effect is not entirely due to insufficient training data. Because the EA method has no training step, its similarly shaped performance curve indicates that larger calibration sets might include residual functions that are irrelevant for future prediction. Again, this result is somewhat expected because the distributions of TAS, TMAX, and PR evolve over time. By including residuals too far in the past, our projection set might contain information that is irrelevant or even counter to future projections. Thus, even if sample size was not a factor, for non-stationary climatic variables, we may still need to limit the size or time period of the calibration set.
@@ 5.2 Forced response in the tails
Given the positive results in Section 5.1, confirming that CE provides reliable UQ for reanalysis data, we use our method to project TAS and PR fields over the next 30 years (Apr. 2024 - Dec. 2054). We use the GP model as our ensemble analysis function, trained on Jan. 1940 - Nov. 2015 data for TAS and Jan 1940 - Nov. 1990 data for PR, and construct CEs on Jul. 2007 - Mar. 2024 data for TAS and Dec 1990 - Mar. 2024 data for PR. Unlike in Section 5.1, we can use the full calibration set to construct each CE. Using the GP, we make three projections over all months between 2024 and 2054 using the IMV, IMV(BC), and the CE to quantify uncertainty and denote these combinations as GP + IMV, GP + IMV(BC), and GP + CE, respectively.